OpenStudy (anonymous):

When a polynomial f(x) is divided by x-1, the reminder is R. The reminder is also R when f(x+1) is divided by ... ? A. x-1 B. x+1 C. x-2 D. x+2 E. x @Callisto I have calculated and got the answer E, but it's wrong...

Parth (parthkohli):

(Deleted my reply because that approach was returning the answer as E... :|)

OpenStudy (anonymous):

@ParthKohli the answer is D but we calculated E

OpenStudy (callisto):

So far I've got: f(x) = (x-1) Q(x) + R f(x+1) = [(x+1)-1]Q(x) + R = x Q(x) + R f(0) = R So, E should be right.

OpenStudy (anonymous):

@Callisto i got the same steps with you

OpenStudy (callisto):

Maybe you need to ask your teacher. I think he/she has made a mistake...

OpenStudy (anonymous):

ok i will ask him tmr. thanks a lot

OpenStudy (callisto):

Do tell us if we're right or not, thanks! :)

OpenStudy (anonymous):

ok i will tell you tmr night here!:) @Callisto @ParthKohli

OpenStudy (anonymous):

the answer is E! sorry for late reply......